# Effective Polaron Dynamics of an Impurity Particle Interacting with a Fermi Gas

### v1

@@ 2.4 Almost-bosonic coherent state
Let $\eta,\xi\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$ and $W_{\lambda}(\eta)\coloneqq e^{\lambda B}\coloneqq e^{\lambda c^{\ast}(\eta)-\lambda c(\eta)}$ for all $\lambda\in[0,1]$, then it holds
$W_{\lambda}(\eta)^{\ast}c(\xi)W_{\lambda}(\eta)$ $=c(\xi)+\lambda\langle\xi,\eta\rangle+\langle\xi,\mathcal{R}^{\lambda}\rangle_{k,\alpha},$
$W_{\lambda}(\eta)^{\ast}c^{\ast}(\xi)W_{\lambda}(\eta)$ $=c^{\ast}(\xi)+\lambda\langle\eta,\xi\rangle+\langle\mathcal{R}^{\lambda},\xi\rangle_{k,\alpha},$
with the short-hand notation
$\langle\xi,\mathcal{R}^{\lambda}\rangle_{k,\alpha}\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\xi_{\alpha}(k)^{\ast}\mathcal{R}^{\lambda}_{\alpha}(k)$ (2.28)
and a $\lambda$-dependent error term $\mathcal{R}^{\lambda}_{\alpha}(k)\coloneqq\int_{0}^{\lambda}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\tau B}$ with $\mathcal{R}\coloneqq\mathcal{R}^{1}$.
@@
We will later give an estimate for $\mathcal{R}$ to show that this term corresponds indeed to a small error. Note that it holds $\mathcal{E}_{\alpha}(k,l)=\mathcal{E}_{\alpha}(l,k)^{\ast}$ for all $l,k\in\Gamma^{\textnormal{nor}}$ and $\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}$ from Lemma A.3 and therefore
$\mathcal{R}^{\lambda}_{\gamma}(l)^{\ast}\xi_{\gamma}(l)$ $=\int_{0}^{\lambda}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{k\in\Gamma^{\textnormal{nor}}}\mathcal{E}_{\gamma}(l,k)\overline{\eta_{\gamma}(k)}\xi_{\gamma}(l)\right)e^{\tau B},$ (2.29)
$\overline{\xi_{\gamma}(l)}\mathcal{R}^{\lambda}_{\gamma}(l)$ $=\int_{0}^{\lambda}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{k\in\Gamma^{\textnormal{nor}}}\overline{\xi_{\gamma}(l)}\eta_{\gamma}(k)\mathcal{E}_{\gamma}(k,l)\right)e^{\tau B}.$ (2.30)
@@
$\langle\eta,\mathcal{R}^{\lambda}\rangle_{k,\alpha}=\langle\mathcal{R}^{\lambda},\eta\rangle_{k,\alpha}$ (2.31)
from which it follows immediately that $\big{(}c^{\ast}(\eta)-c(\eta)\big{)}W_{\lambda}(\eta)=W_{\lambda}(\eta)\big{(}c^{\ast}(\eta)-c(\eta)\big{)}$, i.e. $[B,W_{\lambda}(\eta)]=0$.
@@
The desired statement follows from Duhamel’s formula with the CCR as stated in (2.22)
$[c_{\gamma}(l),e^{\lambda B}]$ $=e^{\lambda B}\int_{0}^{\lambda}\textnormal{d}\tau\ e^{-\tau B}[c_{\gamma}(l),B]e^{\tau B}$
$=e^{\lambda B}\int_{0}^{\lambda}\textnormal{d}\tau\ e^{-\tau B}\left(\eta_{\gamma}(l)+\sum_{k\in\Gamma^{\textnormal{nor}}}\eta_{\gamma}(k)\mathcal{E}_{\gamma}(k,l)\right)e^{\tau B}$
$\equiv\lambda\eta_{\gamma}(l)e^{\lambda B}+e^{\lambda B}\mathcal{R}^{\lambda}_{\gamma}(l)$ (2.32)
Since $c(\xi)\equiv\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\xi_{\alpha}(k)}c_{\alpha}(k)$ is linear the result follows from the above identity. ∎
The following statement shows that $W(\eta)\phi\otimes\Omega$ corresponds to a distribution to the Fock space with expectation approximately being $2\|\eta\|^{2}$.
@@
Let $\psi=\phi\otimes\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$, then it holds for all $\eta\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$
$\langle W(\eta)\psi,\mathcal{N}W(\eta)\psi\rangle=2\|\eta\|^{2}+4\int_{0}^{1}\textnormal{d}\lambda\langle\zeta,\langle\eta,\mathcal{R}^{\lambda}\rangle_{k,\alpha}\zeta\rangle.$
@@
where we used $[\mathcal{N},c_{\alpha}^{\ast}(k)]=2c_{\alpha}^{\ast}(k)$. Therefore
@@
$=2\langle\zeta,B\zeta\rangle+2\|\eta\|^{2}+4\langle\zeta,c(\eta)\zeta\rangle+4\int_{0}^{1}\textnormal{d}\tau\langle\zeta,\langle\eta,\mathcal{R}^{\tau}\rangle_{k,\alpha}\zeta\rangle$ (2.35)
@@
$=\tau\|\eta\|^{2}+\langle\zeta,\langle\eta,\mathcal{R}\rangle_{k,\alpha}\zeta\rangle+\langle\zeta,c(\eta)\zeta\rangle.$ (2.36)
@@
$\langle W(\eta)\zeta,\mathcal{N}W(\eta)\zeta\rangle=2\|\eta\|^{2}+2\langle\zeta,\left(c^{\ast}(\eta)+c(\eta)\right)\zeta\rangle+\langle\zeta,\mathcal{N}\zeta\rangle+4\int_{0}^{1}\textnormal{d}\tau\langle\zeta,\langle\eta,\mathcal{R}^{\tau}\rangle_{k,\alpha}\zeta\rangle.$ (2.37)
The desired result holds for $\zeta\equiv\phi\otimes\Omega$ since $c(\eta)\phi\otimes\Omega=0$.
@@
Let $\eta\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$, there exists a constant $C>0$ such that it holds for all $\tau\in[-1,1],n\in\mathbb{N}$ and $\zeta\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$
@@
$[\mathcal{N},B]=[\mathcal{N},c^{\ast}(\eta)-c(\eta)]=2\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\eta_{\alpha}(k)\big{(}c_{\alpha}^{\ast}(k)+c_{\alpha}(k)\big{)}$ (2.38)
@@
We will later consider time-dependent functions $\eta_{t}$ which are differentiable in $t$ with $\eta_{t},\dot{\eta_{t}}\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$ for each $t\in\mathbb{R}$. In this context, we are interested in the time derivative of the almost-bosonic Weyl operator $W(\eta_{t})$ which can be calculated by the following statement:
@@
Let $\eta_{t}\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$ be differentiable in $t$ with derivative $\dot{\eta_{t}}\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$ for all $t\in\mathbb{R}$. Then it holds for all $t\in\mathbb{R}$
$\partial_{t}W(\eta_{t})=\left(c^{\ast}(\dot{\eta}_{t})-c(\dot{\eta}_{t})+i\textnormal{Im}\langle\dot{\eta}_{t},\eta_{t}\rangle\right)W(\eta_{t})+2i\int_{0}^{1}\textnormal{d}\tau\ W_{(1-\tau)}(\eta_{t})\textnormal{Im}\langle\dot{\eta}_{t},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{t})$
with the shorthand notation $\textnormal{Im}\langle A,B\rangle_{k,\alpha}\coloneqq-\frac{i}{2}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}(A_{\alpha}^{\ast}(k)B_{\alpha}(k)-B_{\alpha}^{\ast}(k)A_{\alpha}(k))$
@@
$=\int_{0}^{1}\textnormal{d}\tau\left\{-B_{s}e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}+e^{-\tau B_{s}}\partial_{s}\partial_{\tau}e^{\tau B_{s}}\right\}$ (2.39)
$=\int_{0}^{1}\textnormal{d}\tau\left\{-B_{s}e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}+e^{-\tau B_{s}}\partial_{s}\left(B_{s}e^{\tau B_{s}}\right)\right\}$ (2.40)
$=\int_{0}^{1}\textnormal{d}\tau\left\{-B_{s}e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}+e^{-\tau B_{s}}\left(\partial_{s}B_{s}\right)e^{\tau B_{s}}+e^{-\tau B_{s}}B_{s}\partial_{s}e^{\tau B_{s}}\right\}$ (2.41)
@@
=$=\left\{[W_{(1-\tau)}(\eta_{s}),c^{\ast}(\dot{\eta}_{s})]-[W_{(1-\tau)}(\eta_{s}),c(\dot{\eta}_{s})]\right\}W_{\tau}(\eta_{s})$ (2.46)
$=W_{(1-\tau)}(\eta_{s})\bigg{\{}(1-\tau)\langle\dot{\eta}_{s},\eta_{s}\rangle-(1-\tau)\langle\eta_{s},\dot{\eta}_{s}\rangle$ (2.47)
@@
$=W(\eta_{s})(1-\tau)2i\text{Im}\langle\dot{\eta}_{s},\eta_{s}\rangle+2iW_{(1-\tau)}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s}).$ (2.49)

### v2

@@ 3.2 Effective coherent state
Let $\eta,\xi\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ and $W_{\sigma}(\eta)\coloneqq e^{\sigma B}\coloneqq e^{\sigma c^{\ast}(\eta)-\sigma c(\eta)}$ for all $\sigma\in[0,1]$, then it holds
$W_{\sigma}(\eta)^{\ast}c(\xi)W_{\sigma}(\eta)$ $=c(\xi)+\sigma\langle\xi,\eta\rangle+\langle\xi,\mathcal{R}^{\sigma}\rangle_{\Gamma},$
$W_{\sigma}(\eta)^{\ast}c^{\ast}(\xi)W_{\sigma}(\eta)$ $=c^{\ast}(\xi)+\sigma\langle\eta,\xi\rangle+\langle\mathcal{R}^{\sigma},\xi\rangle_{\Gamma},$
with $\langle\cdot,\cdot\rangle_{\Gamma}:\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})\times\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})\to\mathbb{C}$ given by
$\langle\xi,\mathcal{R}^{\sigma}\rangle_{\Gamma}\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\xi_{\alpha}(k)^{\ast}\mathcal{R}^{\sigma}_{\alpha}(k)$ (5.1)
and a $\sigma$-dependent error term $\mathcal{R}^{\sigma}_{\alpha}(k)\coloneqq\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\tau B}$.
@@
We will give an estimate for $\mathcal{R}$ to show that this term corresponds indeed to a small error. Note that it holds $\mathcal{E}_{\alpha}(k,l)=\mathcal{E}_{\alpha}(l,k)^{\ast}$ for all $l,k\in\Gamma$ and $\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}$ from Lemma A.3 and therefore
$\mathcal{R}^{\sigma}_{\gamma}(l)^{\ast}\xi_{\gamma}(l)$ $=\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{k\in\Gamma}\mathcal{E}_{\gamma}(l,k)\overline{\eta_{\gamma}(k)}\xi_{\gamma}(l)\right)e^{\tau B},$ (5.2)
$\overline{\xi_{\gamma}(l)}\mathcal{R}^{\sigma}_{\gamma}(l)$ $=\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}\left(\sum_{k\in\Gamma}\overline{\xi_{\gamma}(l)}\eta_{\gamma}(k)\mathcal{E}_{\gamma}(k,l)\right)e^{\tau B}.$ (5.3)
@@
$\langle\eta,\mathcal{R}^{\sigma}\rangle_{\Gamma}=\langle\mathcal{R}^{\sigma},\eta\rangle_{\Gamma}$ (5.4)
from which it follows immediately that $\big{(}c^{\ast}(\eta)-c(\eta)\big{)}W_{\sigma}(\eta)=W_{\sigma}(\eta)\big{(}c^{\ast}(\eta)-c(\eta)\big{)}$, i.e. $[B,W_{\sigma}(\eta)]=0$.
@@
We observe that $W_{\sigma}(\eta)=e^{\sigma B}$ defines a strongly continuous one-parameter semigroup. Thus we can define a derivative and make use of Duhamel’s formula of the form
$e^{-\sigma B}c_{\gamma}(l)e^{\sigma B}=c_{\gamma}(l)+\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}[c_{\gamma}(l),B]e^{\tau B}.\$ (5.5)
The interested reader is referred to [Paz83, EN06] where definitions and properties of operator derivatives are discussed. The desired statement follows with the CCR as stated in (2.22)
$[c_{\gamma}(l),e^{\sigma B}]$ $=e^{\sigma B}\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}[c_{\gamma}(l),B]e^{\tau B}$
$=e^{\sigma B}\int_{0}^{\sigma}\textnormal{d}\tau\ e^{-\tau B}\left(\eta_{\gamma}(l)+\sum_{k\in\Gamma}\eta_{\gamma}(k)\mathcal{E}_{\gamma}(k,l)\right)e^{\tau B}$
$=\sigma\eta_{\gamma}(l)e^{\sigma B}+e^{\sigma B}\mathcal{R}^{\sigma}_{\gamma}(l)$ (5.6)
Since $c(\xi)\equiv\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\xi_{\alpha}(k)}c_{\alpha}(k)$ is linear the result follows from the above identity. ∎
The following statement shows that the number of particles in the state $W(\eta)\phi\otimes\Omega$ corresponds to a random variable with expectation approximately being $2\|\eta\|^{2}$.
@@
Let $\zeta=\phi\otimes\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$, then it holds for all $\eta\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$
$\langle W(\eta)\zeta,\mathcal{N}W(\eta)\zeta\rangle=2\|\eta\|^{2}+4\int_{0}^{1}\textnormal{d}\sigma\langle\zeta,\langle\eta,\mathcal{R}^{\sigma}\rangle_{\Gamma}\zeta\rangle.$
@@
where we used (2.24). Therefore
@@
$=2\langle\zeta,B\zeta\rangle+2\|\eta\|^{2}+4\langle\zeta,c(\eta)\zeta\rangle+4\int_{0}^{1}\textnormal{d}\tau\langle\zeta,\langle\eta,\mathcal{R}^{\tau}\rangle_{\Gamma}\zeta\rangle$ (5.9)
@@
$=\tau\|\eta\|^{2}+\langle\zeta,\langle\eta,\mathcal{R}^{1}\rangle_{\Gamma}\zeta\rangle+\langle\zeta,c(\eta)\zeta\rangle.$ (5.10)
@@
$\langle W(\eta)\zeta,\mathcal{N}W(\eta)\zeta\rangle=2\|\eta\|^{2}+2\langle\zeta,\left(c^{\ast}(\eta)+c(\eta)\right)\zeta\rangle+\langle\zeta,\mathcal{N}\zeta\rangle+4\int_{0}^{1}\textnormal{d}\tau\langle\zeta,\langle\eta,\mathcal{R}^{\tau}\rangle_{\Gamma}\zeta\rangle.$ (5.11)
The desired result holds for $\zeta=\phi\otimes\Omega$ since $c(\eta)\phi\otimes\Omega=0$.
@@
Let $\eta\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$. There exists a constant $C>0$ such that it holds for all $\tau\in[-1,1],n\in\mathbb{N}$ and $\zeta\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$
@@
$[\mathcal{N},B]=[\mathcal{N},c^{\ast}(\eta)-c(\eta)]=2\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\eta_{\alpha}(k)\big{(}c_{\alpha}^{\ast}(k)+c_{\alpha}(k)\big{)}$ (5.12)
@@
Let $\eta_{t}\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ be differentiable in $t$ with derivative $\dot{\eta_{t}}\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ for all $t\in\mathbb{R}$. Then it holds for all $t\in\mathbb{R}$
$\partial_{t}W(\eta_{t})=\left(c^{\ast}(\dot{\eta}_{t})-c(\dot{\eta}_{t})+i\textnormal{Im}\langle\dot{\eta}_{t},\eta_{t}\rangle\right)W(\eta_{t})+2i\int_{0}^{1}\textnormal{d}\tau\ W_{(1-\tau)}(\eta_{t})\textnormal{Im}\langle\dot{\eta}_{t},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{t})$
with the shorthand notation $\textnormal{Im}\langle A,B\rangle_{\Gamma}\coloneqq-\frac{i}{2}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}(A_{\alpha}^{\ast}(k)B_{\alpha}(k)-B_{\alpha}^{\ast}(k)A_{\alpha}(k))$
@@
$=\int_{0}^{1}\textnormal{d}\tau\left(-B_{s}e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}+e^{-\tau B_{s}}\partial_{s}\partial_{\tau}e^{\tau B_{s}}\right)$ (5.13)
$=\int_{0}^{1}\textnormal{d}\tau\left(-B_{s}e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}+e^{-\tau B_{s}}\partial_{s}\left(B_{s}e^{\tau B_{s}}\right)\right)$ (5.14)
$=\int_{0}^{1}\textnormal{d}\tau\left(-B_{s}e^{-\tau B_{s}}\partial_{s}e^{\tau B_{s}}+e^{-\tau B_{s}}\left(\partial_{s}B_{s}\right)e^{\tau B_{s}}+e^{-\tau B_{s}}B_{s}\partial_{s}e^{\tau B_{s}}\right)$ (5.15)
@@
$=\left([W_{(1-\tau)}(\eta_{s}),c^{\ast}(\dot{\eta}_{s})]-[W_{(1-\tau)}(\eta_{s}),c(\dot{\eta}_{s})]\right)W_{\tau}(\eta_{s})$ (5.20)
$=W_{(1-\tau)}(\eta_{s})\bigg{(}(1-\tau)\langle\dot{\eta}_{s},\eta_{s}\rangle-(1-\tau)\langle\eta_{s},\dot{\eta}_{s}\rangle$ (5.21)
@@
$=W(\eta_{s})(1-\tau)2i\text{Im}\langle\dot{\eta}_{s},\eta_{s}\rangle+2iW_{(1-\tau)}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s}).$ (5.23)